Belmont Records produces DVD records. The fixed costs for producing a record are $160,000, and unit cost amount to 85 cents per record. The revenue is $5 per DVD.
(a) How many DVDs must be sold in order to break even?
(b) Find the profit or loss if 20,000 DVD's are sold.
(c) Find the profit or loss if 50,000 DVD's are sold.

Answer by jorel1380(3719)

160000+.85n=5n
160000=4.15n
n=160000/4.15=38554.22
You need to sell 38,554.22 units to break even.
---------------------
20000x5=100000
160000+(20,000x.85)=160000+17000=177000
100000-177000=-77000
Loss of 77000
--------------------------
50000x5=250000
160000+(50000x.85)=160000+42500=202500
250000-202500=47500
Profit of 47500..